OpenStudy (anonymous):

Find the coordinates of the other endpoint when you are given the midpoint and one of the endpoints. Answer in terms of a decimal rounded to the nearest tenth or an improper fraction in simplest form. P1 M P2 (3, 5) (-2, 0) ( , )

OpenStudy (anonymous):

i know i think its -5 , -7 butt show me howwww

OpenStudy (anonymous):

let (x1, y1) be P2. so you'll need to solve these for the coordinates: -2 = (3+x1)/2 0 = (5+y1)/2 in each, solve for the variable to get x1, y1 and that'll be the coordinates of P2

OpenStudy (anonymous):

thanks so much!